OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

subsitution

OpenStudy (texaschic101):

-2x + y = 1 y = 2x + 1 now sub 2x + 1 in for y in the other equation -4x + y = -1 -4x + 2x + 1 = -1 -2x = -1 - 1 -2x = -2 x = 1 now sub 1 in for x in either equation to find y -2x + y = 1 -2(1) + y = 1 -2 + y = 1 y = 2 + 1 y = 3 check... -4x + y = -1 -4(1) + 3 = -1 -4 + 3 = -1 -1 = -1 (correct) x = 1 and y = 3 any questions ?

OpenStudy (anonymous):

holy crap... nope, algebra 1 is the only math i had a problem with. thank you

OpenStudy (texaschic101):

anytime :) algebra 1 is my fav